OpenStudy (anonymous):

What is the solution to the equation? 3x – 6 = –2x + 9 i got 15 is it right

OpenStudy (anonymous):

You know you can check by plugging 15 back into the equation, right?

OpenStudy (anonymous):

BTW, 15 will not be correct, If you were to show what steps you did, I can help you identify the mistake.

OpenStudy (anonymous):

5x-15=0 x=3

OpenStudy (cwrw238):

add 2x to both sides of the equation:- 3x + 2x - 6 = -2x + 2x + 9 5x - 6 = 9 add 6 to both sides 5x - 6 + 6 = 9 + 6 5x = 15 x = 3

OpenStudy (anonymous):

this one..I can help you... 3x – 6 = –2x + 9 transfer to other side so 3x+2x=6+9 then it is 5x=15 therefore x=3 got it?